OpenStudy (candy13106):

a carpenter has a wooden dowel that is 72 cm long. she wants to cut it into 2 pieces so that one is 5 times as long as the other. what are the lentghs of the two pieces?

OpenStudy (solomonzelman):

well, lets call these pieces a and b. So, we get: 72 = a + b Now, we want one piece be 5 times longer than the other, that is: a=5b ------------------------------- your system of equations is: 72=a+b a=5b solve for a and b pliz.

OpenStudy (candy13106):

36

OpenStudy (candy13106):

@SolomonZelman

OpenStudy (anonymous):

yes you are right

OpenStudy (candy13106):

ok r u sure

OpenStudy (anonymous):

im positive lol look at my serious face -_- lol

OpenStudy (solomonzelman):

36 for what?

OpenStudy (candy13106):

a and b

OpenStudy (solomonzelman):

a+b=36?

OpenStudy (candy13106):

yea idk

OpenStudy (anonymous):

how is it not 36 my calculation were on point

OpenStudy (solomonzelman):

You needed to find a and b (separately). I helped you model a system of equations to do this. 72=a+b a=5b

OpenStudy (solomonzelman):

can you tell me the value of a individually, and the value of b individually?

OpenStudy (candy13106):

its 36 for a seperatly and 36 for b